# MsgABD: EveryShipout initializing macros

### v1

@@ Appendix C Test-time Monte-Carlo Averaging
$\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}\leq\mathbb{E}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]$ (C.2)
Therefore, the discussion in the previous subsections extends to the MC-averaged representations as well. Although tighter bounds may be derived for this setting, we will leave that for future works.
Appendix D Additional Figures
In this section, we present some additional figures that demonstrate the negative effects of random edge-dropping, particularly focusing on providing empirical evidence for scenarios not covered by the theory in Section 3. Additionally, we provide visual evidence of the negative effects of DropNode, despite the fact that it preserves sensitivity between nodes (in expectation).
@@ D.1 Symmetrically Normalized Propagation Matrix
The results in Section 3.1 correspond to the use of $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$ for aggregating messages – in each message passing step, only the in-degree of node $i$ is used to compute the aggregation weights of the incoming messages. In practice, however, it is more common to use the symmetrically normalized propagation matrix, $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$, which ensures that nodes with high degree do not dominate the information flow in the graph Kipf & Welling (2017). As in Equation 3.2, we are looking for
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\prod_{l=1}^{L}\hat{\bm{A}}^{\left(\ell\right)}\right]=\ddot{\bm{P}}^{L}$
where $\ddot{\bm{P}}\coloneqq\mathbb{E}_{\mathsf{DE}}[\hat{\bm{A}}^{\text{sym}}]$. While $\ddot{\bm{P}}$ is analytically intractable, we can approximate it using Monte-Carlo sampling. Accordingly, we use 20 samples of $\mathbf{M}$ to compute an approximation of $\ddot{\bm{P}}$, and plot out the entries of $\ddot{\bm{P}}^{L}$, as we did for $\dot{\bm{P}}^{L}$ in Figure 1. The results are presented in Figure 4, which shows that while the sensitivity between nearby nodes is affected to a lesser extent compared to those observed in Figure 1, that between far-off nodes is significantly reduced, same as earlier.
@@ D.2 Upper Bound on Expected Sensitivity
Figure 5: Entries of $\sum_{l=0}^{6}\dot{\bm{P}}^{l}$, averaged after binning node-pairs by their shortest distance.
@@
Black et al. (2023) showed that the sensitivity between any two nodes in a graph can be bounded using the sum of the powers of the propagation matrix. In Section 3.2, we extended this bound to random edge-dropping methods with independent edge masks smapled in each layer:
@@
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]=\zeta_{3}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\mathbb{E}\left[\hat{\bm{A}}\right]^{\ell}\right)_{ij}$
@@
Although this bound does not have a closed form, we can use real-world graphs to study its entries. We randomly sample 100 molecular graphs from the Proteins dataset Dobson & Doig (2003) and plot the entries of $\sum_{l=0}^{6}\dot{\bm{P}}^{\ell}$ (corresponding to DropEdge) against the shortest distance between node-pairs. The results are presented in Figure 5. We observe a polynomial decline in sensitivity as the DropEdge probability increases, suggesting that it is unsuitable for capturing LRIs.
D.3 Test Accuracy versus DropNode Probability
(a) Homophilic datasets
(b) Heterophilic datasets

### v2

@@ C.2 Test-time Monte-Carlo Dropout
$\left\|\frac{\partial}{\partial\bm{x}_{j}}\mathbb{E}\left[\bm{z}^{\left(L\right)}_{i}\right]\right\|_{1}\leq\mathbb{E}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]\leq\ldots$ (C.11)
Therefore, the upper bound results in Section C.1 trivially extend to the MC-averaged representations. Although tighter bounds may be derived for this setting, we leave that for future works.
Appendix D Empirical Sensitivity Analysis
In this section, we present some supplemental figures demonstrating the negative effects of random edge-dropping, particularly focusing on scenarios not covered by the theory. We also elaborate on the setup used for the empirical sensitivity analysis in Section 3.
@@ D.1 Symmetrically Normalized Propagation Matrix
The results in Section 3 correspond to the use of $\hat{\bm{A}}=\hat{\bm{A}}^{\text{asym}}$ for aggregating messages – in each message passing step, only the in-degree of node $i$ is used to compute the aggregation weights of the incoming messages. In practice, however, it is more common to use the symmetrically normalized propagation matrix, $\hat{\bm{A}}=\hat{\bm{A}}^{\text{sym}}$, which ensures that nodes with high out-degree do not dominate the information flow in the graph [47]. As in Equation 3.2, we are looking for
$\ddot{\bm{P}}^{L}\coloneqq\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\prod_{\ell=1}^{L}\hat{\bm{A}}^{\left(\ell\right)}\right]$ (D.1)
where $\ddot{\bm{P}}\coloneqq\mathbb{E}_{\mathsf{DE}}[\hat{\bm{A}}^{\text{sym}}]$. While $\ddot{\bm{P}}$ is analytically intractable, we can approximate it using Monte-Carlo sampling. Accordingly, we use the Cora dataset, and sample 20 DropEdge masks to compute an approximation of $\ddot{\bm{P}}$, and plot out the entries of $\ddot{\bm{P}}^{L}$, as we did for $\dot{\bm{P}}^{L}$ in 1(a). The results are presented in Figure 4, which shows that while the sensitivity between nodes up to 3 hops away is increased, that between nodes farther off is significantly reduced, same as in 1(a).
@@ D.2 Upper Bound on Expected Sensitivity
Figure 5: Entries of $\sum_{\ell=0}^{6}\dot{\bm{P}}^{\ell}$, averaged after binning node-pairs by their shortest distance.
@@
Black et al. [8] showed that the sensitivity between any two nodes in a graph can be bounded using the sum of the powers of the propagation matrix. In Section C.1, we extended this bound to random edge-dropping methods with independent edge masks sampled in each layer:
@@
$\mathbb{E}_{\mathbf{M}^{\left(1\right)},\ldots,\mathbf{M}^{\left(L\right)}}\left[\left\|\frac{\partial\bm{z}^{\left(L\right)}_{i}}{\partial\bm{x}_{j}}\right\|_{1}\right]\leq\zeta_{3}^{\left(L\right)}\left(\sum_{\ell=0}^{L}\mathbb{E}\left[\hat{\bm{A}}\right]^{\ell}\right)_{ij}$
@@
Although this bound does not have a closed form, we can again use the Cora network to study its entries. We plot the entries of $\sum_{\ell=0}^{6}\dot{\bm{P}}^{\ell}$, corresponding to DropEdge, against the shortest distance between node-pairs. The results are presented in Figure 5. We observe an exponential decrease in the sensitivity bound as the distance between nodes increases, suggesting that DropEdge is not suitable for capturing LRIs.
D.3 MC-Approximation of Sensitivity in Nonlinear MPNNs
Given a target node from the Cora dataset [57], we computed the sensitivity of its representation to source nodes up to $L=6$ hops away in ReLU-GCNs of width 32. The raw sensitivities were normalized to obtain influence scores [84]. This was repeated for 25 target nodes, and 25 model--dropout samples were used for each of them. The source nodes were binned by the shortest distance from the corresponding target node, and the influence scores were averaged over each bin to obtain an average influence from nodes $\ell$-hops away.
Why influence scores?
Appendix E Experiments Details
In this section, we expand on the details of the experiments in Section 5. All experiments were run on a server equipped with an Intel(R) Xeon(R) E5-2620 v3 CPU, 62 GB of RAM, 4 $\times$ NVIDIA GeForce GTX TITAN X GPU (12 GB VRAM each), and CUDA version 12.4.
E.1 Descriptions of the Datasets
[…]
Table 4: Statistics of node-classification datasets. Homophily measures from Lim et al. [52].
Dataset Nodes Edges Features Classes Homophily
Homophilic Networks
Cora 2,708 10,556 1,433 7 0.766
CiteSeer 3,327 9,104 3,703 6 0.627
PubMed 19,717 88,648 500 3 0.664
Heterophilic Networks
Chameleon 2,277 36,051 2,325 5 0.062
Squirrel 5,201 216,933 2,089 5 0.025
Actor 7,600 29,926 931 5 0.011
TwitchDE 9,498 306,276 128 2 0.142
Node-classification Tasks. Cora [57], CiteSeer [31] and PubMed [62] are citation networks – their nodes represent scientific publications and an edge between two nodes indicates that one of them has cited the other. The features of each publication are represented by a binary vector, where each index indicates whether a specific word from a dictionary is present or absent. Several studies have showed that these datasets have high homophily in node labels [52, 98] and that they are modelled much better by shallower networks than by deeper ones [95]. Chameleon and Squirrel [69] are networks of English Wikipedia web pages on the respective topics, and the edges between web pages indicate links between them. The task is to predict the average-monthly traffic on each of the web pages. The Actor dataset is induced from a larger film-director-actor-writer network [66]. It is a network of film actors, with edges between those that occur on the same Wikipedia page, and node features are binary vectors denoting the presence of specific keywords in the corresponding Wikipedia entries. The task is to classify actors into five categories based on the content of their Wikipedia pages. Finally, TwitchDE [69] is a network of Twitch users in Germany, with the edges between them representing their mutual follower relationships. The node features are embeddings of the games played by the users, and the task is to predict whether the users use explicit language.
Table 5: Statistics of graph-classification datasets.
Dataset Graphs Nodes Edge Features Classes
Mutag 188  17.9  39.6 7 2
Proteins 1,113  39.1  145.6 3 2
Enzymes 600  32.6  124.3 3 6
Collab 5,000  74.5  4914.4 0 3
IMDb 1,000  19.8  193.1 0 2
Reddit 2,000  429.6  995.5 0 2
Graph-classification Tasks. Following Black et al. [8], Karhadkar et al. [44], we use the graph-classification datasets introduced in Morris et al. [61], which were hypothesized to be long-range tasks. On one hand we have the molecular datasets Mutag, Proteins and Enzymes, and on another we have the social networks Reddit-Binary, IMDb-Binary and Collab. Their statistics are presented in Table 5.
Mutag [17] consists of nitroaromatic compounds, with the objective of predicting their mutagenicity, i.e. the ability to cause genetic mutations, in cells in Salmonella typhimurium. Each compound is represented as a graph, where nodes correspond to atoms, represented by their type using one-hot encoding, and edges denote the bonds between them. Proteins [19, 9] is a collection of proteins classified as enzymes or not. The molecules are represented as a graph of amino acids, with edges between those separated up to 6Å apart. The Enzymes [9, 73] dataset is represented similarly, and the task is to classify the proteins into one of 6 Enzyme Commission (EC) numbers – a system to classify enzymes based on the reactions they catalyze.
Collab [87] is a scientific collaboration dataset where each graph represents a researcher’s ego network. Nodes correspond to researchers and their collaborators, with edges indicating co-authorship. Each network is labeled based on the researcher’s field, which can be High Energy Physics, Condensed Matter Physics, or Astrophysics. IMDb-Binary [87] is a movie collaboration dataset containing the ego-networks of 1,000 actors and actresses from IMDB. In each graph, nodes represent actors, with edges connecting those who have co-starred in the same film. Each graph is derived from either the Action genre or Romance. Finally, Reddit-Binary [87] comprises graphs representing online discussions on Reddit, where nodes correspond to users and edges indicate interactions through comment responses. Each graph is labeled based on whether it originates from a Q&A or discussion-based subreddit.
For the molecular datasets, we use the node features supplied by PyG [24], but since they are unavailable for the social networks, we set scalar features $\bm{x}_{i}=\left(1\right)$ for all nodes in these datasets, following [44].
E.2 Training Configurations
Dataset Splits. For the SyntheticZINC task, we use the train-val-test splits provided in PyG. For the homophilic (citation) networks, we use the ‘full’ split [12], as provided in PyG, and for the heterophilic networks, we randomly sample 60% of the nodes for training, 16% for validation, and 24% for testing. On the other hand, for the graph classification tasks, we sample 80% of the graphs for training, and 10% each for validation and testing, following [44, 8].
Model Architecture. We standardize most of the hyperparameters across all experiments to isolate the effect of random dropping. Specifically, we use symmetric normalization of the adjacency matrix to compute the edge weights for GCN, and we set the number of attentions heads for GAT to 2 in order to keep the computational load manageable, while at the same time harnessing the expressiveness of the multi-headed self-attention mechanism. For the SyntheticZINC dataset, we fix the size of the hidden representations at 16, while we fix them to 64 for all the real-world datasets. In all settings, a linear transformation is applied to the node features before message-passing. Afterwards, a bias term is added and then the ReLU nonlinearity is applied. Finally, a linear readout layer is used to compute the regressand (for regression tasks) or logits (for classification).
Dropping Probability. For the synthetic datasets, we experiment with a NoDrop baseline, and DropEdge, Dropout and DropMessage, each with $q=0.2$ and $q=0.5$. For the real-world datasets, the dropping probabilities are varied as $q=0.1,0.2,\ldots,0.9$, so as to reliably find the best performing configuration. We adopt the common practice of turning the dropping methods off at test-time ($q=0$), isolating the effects on optimization and generalization, which our theory does not address.
DropSens Configurations. For DropSens, we use 4 possible values for proportion of information preserved over corss-edges, $c=0.5,0.8,0.9,0.95$. Since the dropping probability, $q_{i}$, increases with the in-degree of the target node, $d_{i}$, the proportion of all edges dropped could become very high, especially with large $c$. Therefore, we clip the value of $q_{i}$ by 4 possible choices, $q_{i}\leq q_{\max}\in\left\{0.2,0.3,0.5,0.8\right\}$. We exclude the following configurations: $\left(c,q_{\max}\right)\in\left\{\left(0.5,0.2\right),\left(0.5,0.3\right),\left(0.5,0.5\right),\left(0.8,0.2\right),\left(0.8,0.3\right)\right\}$, since they use the same dropping probability ($=q_{\max}$) for each edge, and are therefore equivalent to DropEdge. In summary, we test with a total of 11 configurations for DropSens to find the best one for each task.
Optimization Algorithm. The models are trained using the Adam optimizer [46]. On the SyntheticZINC dataset, the models are trained with a learning rate of $2\times 10^{-3}$ and a weight decay of $1\times 10^{-4}$, for a total of 200 epochs. On the real-world datasets, we use a learning rate of $1\times 10^{-3}$ and no weight decay, following Karhadkar et al. [44], Black et al. [8]. Here, we cap the maximum number of epochs at 300. In both cases, the learning rate is reduced by a factor of $1\times 10^{-1}$ if the validation loss fails to improve beyond a relative threshold of $1\times 10^{-4}$ for 10 epochs, again following Karhadkar et al. [44], Black et al. [8].
Number of Independent Runs. We perform only 10 independent runs on the SyntheticZINC dataset due to its consistently low variance in performance, as also observed by Giovanni et al. [33]. For real-world datasets, we conduct 20 runs to identify the best-performing dropping configurations. We then perform a one-sided t-test to assess whether dropout improves performance, using a 90% confidence-level ($\alpha=0.1$) and targeting a statistical power of 0.9 ($\beta=0.1$). Under the assumption that the dropping method offers superior performance, detecting a medium effect size of $0.5$ [14] requires approximately 53 samples per group according to standard power analysis. We round this to 50, and accordingly perform 30 additional runs for the final comparison of the best-performing dropping configuration with the NoDrop baseline.
E.3 DropSens Implementation
import warnings
import sympy
from sympy.abc import x as q
import torch
from torch_geometric.utils import degree, contains_self_loops
def drop_sens(
edge_index: torch.Tensor,
c: float,
max_drop_prob: float = None
if max_drop_prob is None:
max_drop_prob = 1.
# Assuming edge index does not have self loops
if contains_self_loops(edge_index):
warnings.warn("Degree computation in DropSens assumes absence",
"of self-loops, but the edge_index passed has them.")
degrees = degree(edge_index[1]).int() # Node index -> in-degree
ds = torch.unique(degrees).tolist() # Sorted array
mapper = torch.nan * torch.ones(ds[-1]+1)
mapper[ds] = max_drop_prob # Node degree -> dropping prob
for d_i in ds:
q_i = float(sympy.N(sympy.real_roots(
(1-c)*d_i*(1-q) - q + q**(d_i+1))[-2] # Following Equation 4.1
)) if d_i > 0 else 0.
if q_i > max_drop_prob:
# Because q monontonic wrt d, and ds is sorted
break
mapper[d_i] = q_i
in_degrees = degrees[edge_index[1]] # Edge index -> in-degree of target node
qs = mapper[in_degrees] # Edge index -> dropping probability
edge_mask = qs <= torch.rand(edge_index.size(1))
edge_index = edge_index[:, edge_mask]
return edge_index, edge_mask
Listing 1: DropSens Implementation
In 1, we present the DropSens implementation used in our experiments, relying mainly on SymPy [59].
Figure 6: Edge-wise dropping probabilities under DropSens for varying values of $c$, along with corresponding approximations as in Equation E.1.
Unfortunately, computing the roots of Equation 4.1 becomes slow when the in-degree $d_{i}$ is large – a common scenario in large networks. This issue is especially pronounced when the proportion of information preserved $c$ is large, as the dropping threshold is only met at a higher value of $d_{i}$. To address this computational challenge, we propose an approximation:
$1-c=\frac{q_{i}-q_{i}^{d_{i}+1}}{d_{i}\left(1-q_{i}\right)}\approx\frac{q_{i}}{d_{i}\left(1-q_{i}\right)}\quad\Longrightarrow\quad q_{i}\approx\frac{\left(1-c\right)d_{i}}{1+\left(1-c\right)d_{i}}$ (E.1)
This approximation becomes increasingly accurate as $c$ increases – since more information needs to be preserved, $q_{i}$ needs to be small, and hence, $q_{i}^{d_{i}+1}\to 0$.
Figure 6 shows the DropSens probabilities for masking incoming messages based on the in-degree of the target nodes, along with the corresponding approximations. It is clear to see that the approximation gets increasingly more accurate for lower proportions of information loss.
Appendix F Supplementary Experiments
F.1 Test Accuracy versus DropEdge Probability
In Section 3, we studied the effect of edge-dropping probability on sensitivity between nodes at different distances. However, this analysis may be insufficient to precisely predict the impact on model performance since DropEdge-variants significantly affect the optimization trajectory as well. To learn more about the relationship between test-time performance and dropping probability, we evaluate DropEdge-GCNs on the heterophilic datasets; the results are shown in Figure 7. Clearly, on Chameleon, Squirrel and TwitchDE, the performance degrades with increasing dropping probability, as was suggested by Theorem 3.1 and 1(a). Surprisingly, the trends are significantly monotonic with GCNs of all depths, $L=2,4,6,8$.
F.2 Remark on DropNode